INFRARED SPECTROSCOPY OF N_{2}O CLUSTERS

M. AFSHARI, M. DEHGHANI, Z. ABUSARA, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, 2500 University Drive NW, Calgary, Alberta T2N 1N4, Canada; \underline{A.R.W. MCKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

Following our observation of the polar dimer of OCS, we searched the region of the N-N stretching fundamental of N_{2}O using a tunable diode laser to probe a pulsed supersonic slit jet. In addition to the relatively strong ν_{3} band of the centrosymmetric lowest energy isomer of N_{2}O dimer, we observed a weak band at around 2226.45 cm-1 which we have assigned to the higher energy polar isomer of N_{2}O dimer. 140 transitions from an a/b type hybrid band were assigned and fitted to a planar asymmetric top structure. Our lower state parameters should enable the observation of the microwave spectrum of polar (N_{2}O)_{2}. A trimer band of N_{2}O has also been assigned based on lower state combination differences from previously published works. In contrast to OCS where we have observed four tetramer bands, only one tetramer band around 2237.4 cm-1 has been found to date. This band is best described as an asymmetric top with an accidental spherical top structure. Isotopic studies of these bands are presently underway.